# Generating Analytic Specifications for Data Visualization from Natural Language Queries using Large Language Models

### v1

@@ Abstract
Recently, large language models (LLMs) have shown great promise in translating natural language (NL) queries into visualizations, but their “black-box” nature often limits explainability and debuggability. In response, we present a comprehensive text prompt that, given a tabular dataset and an NL query about the dataset, generates an analytic specification including (detected) data attributes, (inferred) analytic tasks, and (recommended) visualizations. This specification captures key aspects of the query translation process, affording both explainability and debuggability. For instance, it provides mappings from the detected entities to the corresponding phrases in the input query, as well as the specific visual design principles that determined the visualization recommendations. Moreover, unlike prior LLM-based approaches, our prompt supports conversational interaction and ambiguity detection capabilities. In this paper, we detail the iterative process of curating our prompt, present a preliminary performance evaluation using GPT-4, and discuss the strengths and limitations of LLMs at various stages of query translation. The prompt is open-source and available to view at https://github.com/nl4dv/nl4dv.
@@
In this work, we present a new LLM-based text prompt (NL4DV-LLM) that, like NL4DV [29, 26], returns an analytic specification containing data attributes, analytic tasks, and relevant visualizations.
This specification, presented as a structured JSON object (preferred by developers) or a step-by-step natural language explanation (preferred by users), affords explainability and debuggability by documenting key aspects of the query translation process. For example, it provides the mappings between the detected entities and the corresponding phrases in the input query as well as the specific visual design principles that determined the visualization recommendation. Furthermore, this prompt offers conversational interaction and ambiguity detection functionalities which are currently unsupported in other LLM-based NL2VIS systems. Essentially, users can ask follow-up queries to alter previously generated analytic specification(s) based on their evolving needs. If a query also contains ambiguities (e.g. a query phrase that can map to multiple data attributes), the prompt outputs multiple visualizations, one for each ambiguous entity. However, the prompt does not support query resolution since it is a programmatic capability outside the scope of query translation [26]. Figure 1 highlights this key difference in the capabilities of NL4DV and NL4DV-LLM.
@@
We also conducted a preliminary evaluation of NL4DV-LLM against NL4DV using the NLVCorpus dataset [38] and GPT-4 [33] as the prompt’s engine. We found that in our corpus of 740 queries across three datasets, our prompt achieved an accuracy of 87.02% compared to NL4DV’s accuracy of 64.05%.
However, on average, the LLM took around 25 seconds to generate analytic specifications, which can be a potentially unreasonable wait time for users.
We discuss the tradeoffs and strengths of this prompt, which we hope guides future developments in natural language to visualization interfaces. Our primary contributions are the following:
@@ 1.1.5 Data Subset
For the LLM to detect references to attributes and records in the input query, it needs access to the entire dataset. However, due to a token limit imposed by the GPT-4 API, it is not possible to include the entire dataset as part of the prompt. Consequently, we select all dataset headers (columns) and randomly subset ten records (rows) for the prompt. While this choice alleviates the token-limit concern, it still has two limitations. First, for really wide datasets, the large number of columns may still exceed the token limit; and second, the ten sample rows may not be representative of the entire dataset and may result in false negatives in the query translating process.
To resolve cases where the amount of dataset columns exceeds the prompt’s token limit, users can leverage methods in exploratory data analysis such as Linear Discriminant Analysis [40] that only includes the most relevant columns as part of the prompt.
@@ 1.2 Prompt Iterations
Next, in Version 3, we enhanced the analytic task JSON by introducing support for other analytic tasks, namely Sort, Filter, and Find Extremum.
An Attributes and Visual Encodings Description property was also included for each task, providing the LLM heuristics on task inference for underspecified queries. Version 4 provided additional instructions to handle ambiguous and underspecified queries since earlier prompt versions were unable to generate accurate visualizations for such query types.
@@ 2.1 Setup and Design
We provide the dataset corpus (used for evaluation), text prompt (NL4DV-LLM), annotations on the prompt’s outputs, and a gallery of sample visualization outputs in the supplemental material.
@@ 2.3 Results and Discussion
Out of the 740 queries, NL4DV-LLM generated 644 accurate responses, resulting in an accuracy rate of 87.02%. NL4DV generated 474 accurate responses, resulting in an accuracy rate of 64.05%. The prompt’s accuracy was mostly consistent across the three test datasets: 84.98% on the cars’ dataset, 89.89% accuracy on the movies’ dataset, and 86.5% on the superstore dataset. However, NL4DV’s accuracies on the movies’ dataset (75.49%) and cars dataset (70.73%) were significantly higher than the superstore dataset (40%), potentially due to the superstore dataset’s complexity and high syntactic similarity among its attributes.
A previous evaluation reported that learning-based NL2VIS system ncNet [21] had an accuracy of about 45% and another LLM-based system chartGPT [41] had an accuracy of about 79%, albeit on the nvBench dataset and a subtle distinction in its definition of accuracy [41].
These findings suggest that our prompt’s accuracy is comparable to, if not slightly higher than, previous systems.
Notably, our prompt handled a wide range of query structures and forms, including underspecified and ambiguous queries. For example, our prompt was able to apply transformations and computations to dataset values to create new “derived” attributes (e.g. creating a new attribute Profit from Production Budget and Worldwide Gross in the movies dataset), which semantic parsing-based approaches cannot generally support.
However, there were a few queries that resulted in inaccurate outputs. Common reasons for inaccuracy were malformed response JSON objects or well-formed JSON objects with incorrect Vega-Lite syntax. Such occurrences can undermine user trust, highlighting the need for future systems to implement appropriate safeguards. In addition, certain outputs were misleading due to incorrect associations between the data attributes and the visual encodings.
For example, for the query, “Show total profit across genres,” the y-axis might be labeled “Total profit” but actually use “Worldwide Gross” as the data field, making the visualization confusing.

### v2

@@ Abstract
Recently, large language models (LLMs) have shown great promise in translating natural language (NL) queries into visualizations, but their “black-box” nature often limits explainability and debuggability. In response, we present a comprehensive text prompt that, given a tabular dataset and an NL query about the dataset, generates an analytic specification including (detected) data attributes, (inferred) analytic tasks, and (recommended) visualizations. This specification captures key aspects of the query translation process, affording both explainability and debuggability. For instance, it provides mappings from the detected entities to the corresponding phrases in the input query, as well as the specific visual design principles that determined the visualization recommendations. Moreover, unlike prior LLM-based approaches, our prompt supports conversational interaction and ambiguity detection capabilities. In this paper, we detail the iterative process of curating our prompt, present a preliminary performance evaluation using GPT-4, and discuss the strengths and limitations of LLMs at various stages of query translation. The prompt is open-source and integrated into NL4DV, a popular Python-based natural language toolkit for visualization, which can be accessed at https://nl4dv.github.io.
@@
In this work, we present a new LLM-based text prompt (NL4DV-LLM) that, like NL4DV [29, 26], returns an analytic specification containing data attributes, analytic tasks, and relevant visualizations. This specification, presented as a structured JSON object (preferred by developers) or a step-by-step natural language explanation (preferred by users), affords explainability and debuggability by documenting key aspects of the query translation process. For example, it provides the mappings between the detected entities and the corresponding phrases in the input query as well as the specific visual design principles that determined the visualization recommendation. Furthermore, this prompt offers conversational interaction and ambiguity detection functionalities which are currently unsupported in other LLM-based NL2VIS systems. Essentially, users can ask follow-up queries to alter previously generated analytic specification(s) based on their evolving needs. If a query also contains ambiguities (e.g. a query phrase that can map to multiple data attributes), the prompt outputs multiple visualizations, one for each ambiguous entity. However, the prompt does not support query resolution since it is a programmatic capability outside the scope of query translation [26]. Figure 1 highlights this key difference in the capabilities of NL4DV and NL4DV-LLM.
@@
We also conducted a preliminary evaluation of NL4DV-LLM against NL4DV using the NLVCorpus dataset [38] and GPT-4 [33] as the prompt’s engine. We found that in our corpus of 740 queries across three datasets, our prompt achieved an accuracy of 87.02% compared to NL4DV’s accuracy of 64.05%. However, on average, the LLM took around 25 seconds to generate analytic specifications, which can be a potentially unreasonable wait time for users. We discuss the tradeoffs and strengths of this prompt, which we hope guides future developments in natural language to visualization interfaces. Our primary contributions are the following:
@@ 1.1.5 Data Subset
For the LLM to detect references to attributes and records in the input query, it needs access to the entire dataset. However, due to a token limit imposed by the GPT-4 API, it is not possible to include the entire dataset as part of the prompt. Consequently, we select all dataset headers (columns) and randomly subset ten records (rows) for the prompt. While this choice alleviates the token-limit concern, it still has two limitations. First, for really wide datasets, the large number of columns may still exceed the token limit; and second, the ten sample rows may not be representative of the entire dataset and may result in false negatives in the query translating process. To resolve cases where the amount of dataset columns exceeds the prompt’s token limit, users can leverage methods in exploratory data analysis such as Linear Discriminant Analysis [40] that only includes the most relevant columns as part of the prompt.
@@ 1.2 Prompt Iterations
Next, in Version 3, we enhanced the analytic task JSON by introducing support for other analytic tasks, namely Sort, Filter, and Find Extremum. An Attributes and Visual Encodings Description property was also included for each task, providing the LLM heuristics on task inference for underspecified queries. Version 4 provided additional instructions to handle ambiguous and underspecified queries since earlier prompt versions were unable to generate accurate visualizations for such query types.
@@ 2.1 Setup and Design
We provide the dataset corpus (used for evaluation), text prompt (NL4DV-LLM), annotations on the prompt’s outputs, and a gallery of sample visualization outputs which can be accessed at https://github.com/nl4dv/NL4DV-LLM-supplemental-material.
@@ 2.3 Results and Discussion
Out of the 740 queries, NL4DV-LLM generated 644 accurate responses, resulting in an accuracy rate of 87.02%. NL4DV generated 474 accurate responses, resulting in an accuracy rate of 64.05%. The prompt’s accuracy was mostly consistent across the three test datasets: 84.98% on the cars’ dataset, 89.89% accuracy on the movies’ dataset, and 86.5% on the superstore dataset. However, NL4DV’s accuracies on the movies’ dataset (75.49%) and cars dataset (70.73%) were significantly higher than the superstore dataset (40%), potentially due to the superstore dataset’s complexity and high syntactic similarity among its attributes. A previous evaluation reported that learning-based NL2VIS system ncNet [21] had an accuracy of about 45% and another LLM-based system chartGPT [41] had an accuracy of about 79%, albeit on the nvBench dataset and a subtle distinction in its definition of accuracy [41]. These findings suggest that our prompt’s accuracy is comparable to, if not slightly higher than, previous systems.
Notably, our prompt handled a wide range of query structures and forms, including underspecified and ambiguous queries. For example, our prompt was able to apply transformations and computations to dataset values to create new “derived” attributes (e.g. creating a new attribute Profit from Production Budget and Worldwide Gross in the movies dataset), which semantic parsing-based approaches cannot generally support. However, there were a few queries that resulted in inaccurate outputs. Common reasons for inaccuracy were malformed response JSON objects or well-formed JSON objects with incorrect Vega-Lite syntax. Such occurrences can undermine user trust, highlighting the need for future systems to implement appropriate safeguards. In addition, certain outputs were misleading due to incorrect associations between the data attributes and the visual encodings. For example, for the query, “Show total profit across genres,” the y-axis might be labeled “Total profit” but actually use “Worldwide Gross” as the data field, making the visualization confusing to the user.